### v2

# Yangming Zhang*    Wenqi Jia* Dept. of Computer Science

Abstract
footnotetext: *Yangming Zhang is the leading co-first author, while Wenqi Jia is the secondary co-first author. Miao Yin is the corresponding author.
3D Gaussian Splatting (3DGS) has emerged as a mainstream for novel view synthesis, leveraging continuous aggregations of Gaussian functions to model scene geometry. However, 3DGS suffers from substantial memory requirements to store the multitude of Gaussians, hindering its practicality. To address this challenge, we introduce GaussianSpa, an optimization-based simplification framework for compact and high-quality 3DGS. Specifically, we formulate the simplification as an optimization problem associated with the 3DGS training. Correspondingly, we propose an efficient “optimizing-sparsifying” solution that alternately solves two independent sub-problems, gradually imposing strong sparsity onto the Gaussians in the training process. Our comprehensive evaluations on various datasets show the superiority of GaussianSpa over existing state-of-the-art approaches. Notably, GaussianSpa achieves an average PSNR improvement of 0.9 dB on the real-world Deep Blending dataset with 10$\times$ fewer Gaussians compared to the vanilla 3DGS. Our project page is available at https://noodle-lab.github.io/gaussianspa.
Figure 1: We present GaussianSpa, enabling high-quality and compact view synthesis with superior rendering of details. Compared to the existing state-of-the-art method, Mini-Splatting [20], our GaussianSpa captures detail-rich textures and background more accurately with fewer Gaussians.
1 Introduction
Novel view synthesis has become a pivotal area in computer vision and graphics, driving advancements in applications such as virtual reality, augmented reality, and immersive media experiences [21]. NeRF [35] has recently gained prominence in this domain because it can generate high-quality, photorealistic images from sparse input views by representing scenes as continuous volumetric functions based on neural networks. However, NeRF often requires extensive computational resources and long training times, making it less practical for real-time applications and large-scale reconstructions.
3D Gaussian Splatting (3DGS) [25] has emerged as a powerful alternative, leveraging continuous aggregations of Gaussian functions to model scene geometry and appearance. Unlike NeRF, which relies on neural networks to approximate volumetric radiance fields, 3DGS directly represents scenes using a collection of Gaussians. This approach excels in capturing details and smooth transitions, offering significant advantages in training and rendering speed. 3DGS achieves superior visual fidelity [28] compared to NeRF while reducing computational overhead, making it more suitable for interactive applications that demand quality and performance.
Despite its strengths, 3DGS suffers from significant memory requirements that hinder its practicality. The main issue is the massive memory consumption associated with storing a large number of Gaussians that are needed to represent complex scenes. Each Gaussian occupies memory space for its parameters, including position, covariance, and color attributes. In densely sampled scenes, the sheer volume of Gaussians leads to memory usage that exceeds the capacity of typical hardware, making it challenging to handle higher-resolution scenes and limiting its applicability in resource-constrained environments.
Existing works, e.g., Mini-Splatting [20], LightGaussian [19], LP-3DGS [54], EfficientGS [31], and RadSplat [40], have predominantly focused on mitigating this issue by removing a certain number of Gaussians. Techniques such as pruning and sampling aim to discard unimportant Gaussians based on hand-crafted criteria such as opacity [25, 53], importance score (hit count) [20, 19, 40], dominant primitives [31], and binary mask [54, 29]. However, these criteria are generally utilized to determine the importance of Gaussian points from a single heuristic perspective, limiting robustness in dynamic scenes or under varying lighting conditions. Moreover, the sudden one-shot removal may cause permanent loss of Gaussians that are crucial to visual synthesis, making it challenging to recover the original performance after even long-term training, as shown in Figure 2. Consequently, while these methods can alleviate memory and storage burdens to some extent, they often lead to sub-optimal rendering outcomes with loss of details and visual artifacts, thereby compromising the quality of the synthesized views.
In this paper, we present an optimization-based simplification framework, GaussianSpa, for compact and high-quality Gaussian Splatting. In the proposed framework, we formulate 3DGS simplification as a constrained optimization problem under a target number of Gaussians. Then, we propose an efficient “optimizing-sparsifying" solution for the formulated problem by splitting it into two simple sub-problems that are alternately solved in the “optimizing" step and the “sparsifying" step. Instead of permanently removing a certain number of Gaussians, GaussianSpa incorporates the “optimizing-sparsifying” algorithm into the training process, gradually imposing a substantial sparse property onto the trained Gaussians. Hence, our GaussianSpa can simultaneously enjoy maximum information preservation from the original Gaussians and a desired number of reduced Gaussians, providing compact 3DGS models with high-quality rendering. Overall, our contributions can be summarized as follows:
  • We propose a general 3DGS simplification framework that formulates the simplification objective as an optimization problem and solves it in the 3DGS training process. In solving the formulated optimization problem, our proposed framework gradually restricts Gaussians to the target sparsity constraint without explicitly removing a specific number of points. Hence, GaussianSpa can maximally maintain and smoothly transfer the information to the sparse Gaussians from the original model.
  • We propose an efficient “optimizing-sparsifying" solution for the formulated problem, which can be integrated into the 3DGS training with negligible costs, separately solving two sub-problems. In the “optimizing" step, we optimize the original loss function attached by a regularization with gradient descent. In the “sparsifying" step, we analytically project the auxiliary Gaussians onto the constrained sparse space.
  • We comprehensively evaluate GaussianSpa through extensive experiments on various complex scenes, demonstrating improved rendering quality compared to existing approaches. Particularly, with as high as 10$\times$ fewer number of Gaussians than the vanilla 3DGS, GaussianSpa achieves an average 0.4 dB improvement on the Mip-NeRF 360 [4] and Tanks&Temples [27] datasets, 0.9 dB on Deep Blending [23] dataset. Furthermore, we conduct various visual quality assessments, showing that GaussianSpa exhibits high-quality rendering of details and sparse 3D Gaussian views.
2 Related Work
2.1 Novel View Synthesis
Novel View Synthesis (NVS) [2] is a technique in computer vision and graphics to generate images of a 3D scene from unseen viewpoints based on existing multi-view data. A seminal work in this field is NeRF [35], which uses an implicit neural network to represent a scene as a continuous 5-dimensional function, mapping 3D coordinates and viewing directions to color and density. NeRF introduces positional encoding and a hierarchical sampling strategy to improve the ability of the MLP [6] to represent high-frequency image details and accelerate the rendering process. However, NeRF’s reliance on single narrow-ray sampling poses challenges in addressing aliasing and blur artifacts. To mitigate this, Mip-NeRF [3] represents the scene at a continuous-valued scale, enabling anti-aliased rendering by casting conical frustums from each pixel at various scales. Additionally, Mip360 [4] extends NeRF’s capabilities to unbounded scenes, utilizing a non-linear scene parameterization, online distillation, and a novel distortion-based regularizer. More recently, an explicit point-based method, 3D Gaussian Splatting (3DGS), which leverages continuous aggregations of Gaussian functions to model scene geometry, has emerged as a point-based alternative, demonstrating significant improvements in rendering quality and speed.
Figure 2: PSNR curves of human-crafted criteria-based pruning methods. All methods remove 85% of the Gaussians at iteration 25K.
2.2 Efficient 3DGS
3DGS significantly improves rendering quality and computational efficiency compared to traditional methods such as NeRF [35], which are often resource-intensive and computationally slow. However, the demands of real-time rendering and the reconstruction of unbounded, large-scale 3D scenes underscore the requirement for model efficiency [10]. Optimization for 3DGS focuses on four main categories to effectively reduce model size: Gaussian simplification, vector quantization, spherical harmonics optimization, and hybrid compression.
Gaussian Simplification. Gaussian simplification involves removing Gaussians to compress 3DGS models into compact forms, significantly reducing memory costs. Prior approaches [20, 40, 34, 54, 32, 46, 30, 13, 26] mainly focus on pruning redundant Gaussians according to hand-crafted importance criteria. Mini-Splatting [20] introduces blur splitting and depth reinitialization to address overlapping and under-reconstructed problems, using stochastic sampling to better preserve geometry based on importance scores. Radsplatting [40] computes an importance score by aggregating ray contributions of Gaussians across all input images, replacing summation with a max operator for robustness to varying scene coverage. Taming 3DGS [34] implements steerable densification by calculating scores based on pixel saliency, gradients, and Gaussian attributes, then selectively cloning or splitting high-scoring Gaussians. LP-3DGS [54] applies a trainable binary mask with Gumbel-Sigmoid to automatically optimize pruning, improving efficiency while maintaining quality.
Vector Quantization. Vector quantization [50] aims to compress data using discrete entries from a learned codebook, reducing data redundancy. Several works [22, 38, 12, 51, 42, 36, 39] apply vector quantization for compressing 3DGS models. For instance, EAGLES [22] quantizes attributes like spherical harmonics (SH) coefficients and opacity to create latent representations for per-point compression, using an MLP decoder to reconstruct the original attributes with high fidelity. CompGS [38] applies a K-means-based vector quantization approach to compress 3DGS parameters, significantly reducing storage and rendering time with minimal quality loss. RDO-Gaussian [51] performs end-to-end rate-distortion optimization for 3D Gaussian Splatting, combining dynamic pruning and entropy-constrained vector quantization to reduce storage while maintaining quality.
Spherical Harmonics Optimization. Among 3DGS parameters, spherical harmonics (SH) coefficients take up a significant portion, requiring (45+3) floating-point values per Gaussian for a degree of 3 (as used in [25]), which accounts for around 80% of the total attribute volume. Prior works [19, 29] optimize SH coefficients to reduce storage overhead. For instance, EfficientGS [31] retains zeroth-order SH for all Gaussians, increasing the order when necessary.
Hybrid Compression. In addition to methods focused on individual aspects of 3DGS compression, several recent approaches [19, 31, 29, 52] combine multiple techniques for hybrid compression to achieve further storage reduction. LightGaussian [19] prunes 3DGS points using a global importance score, applies vector quantization to compress SH coefficients, and reduces SH degree through knowledge distillation combined with pseudo-view augmentation. EfficientGS [31] not only keeps zeroth-order SH for all Gaussians, increasing the order only when needed but also classifies Gaussians into steady and non-steady states, densifying non-steady Gaussians while retaining only the top-K most important Gaussians per ray. CompactGaussian [29] introduces a learnable mask to mark redundant Gaussians and residual vector quantization for scale and rotation and replaces SH with a grid-based neural field to represent colors.
Although these methods improve 3DGS efficiency, they lead to irreversible information loss during the Gaussian simplification process. Specifically, those methods perform simplification by a one-shot pruning on the Gaussians without prominent sparse characteristics, causing a notable PSNR degradation that is challenging to recover, as shown in Figure 2. Consequently, when applying additional compressions like SH optimization and vector quantization after simplification, the compressed 3DGS model often exhibits further performance loss due to the accumulated errors. Our work focuses on Gaussian simplification by formulating it as an optimization problem and imposing substantial sparsity onto Gaussians, mitigating performance issues with improved rendering quality at the first stage to achieve efficient 3DGS.
Figure 3: Overall workflow of our proposed GaussianSpa framework.
3 Methodology: GaussianSpa
3.1 Background of 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) explicitly represents scenes using a collection of point-based 3D continuous Gaussians. Specifically, each Gaussian $G$ is characterized by its covariance matrix $\bm{\Sigma}$ and center position $\bm{\mu}$ as
$G(\bm{x})=\exp\left(-\frac{1}{2}(\bm{x}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})\right),$ (1)
where $\bm{x}$ is an arbitrary location in the 3D scene. The covariance matrix $\bm{\Sigma}$ is generally composed of $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$ with a rotation matrix $\bm{R}$ and a scale matrix $\bm{S}$, ensuring the positive definite property. Additionally, each Gaussian is associated with an opacity $a$ and spherical harmonics (SH) coefficients.
When rendering a 2D image from the 3D scene at a certain camera angle, 3DGS projects 3D Gaussians to that 2D plane and calculates the projected 2D covariance matrix by
$\bm{\Sigma^{\prime}}=\bm{J}\bm{W}\bm{\Sigma}\bm{W}^{\top}\bm{J}^{\top},$ (2)
where $\bm{W}$ is the view transformation, and $\bm{J}$ is the Jacobian of the affine approximation of the projective transformation [55]. Then, for each pixel on the 2D image, the color $C$ is computed by blending all $N$ ordered Gaussians contributing to the pixel as
$C=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (3)
Here, $c_{i}$ and $\alpha_{i}$ denote the view-dependent color computed from the corresponding SH coefficients and the rendering opacity calculated by the Gaussian opacity $a_{i}$, respectively, corresponding to the $i$-th Gaussian.
In the training process [25], the Gaussians are initialized from a sparse point cloud generated by Structure-from-Motion (SfM) techniques [47]. Then, the number of Gaussians is adjusted by a densification algorithm [25] according to the reconstruction geometry. Afterward, each Gaussian’s variables, including the center position, rotation matrix, scaling matrix, opacity, and SH coefficients, are optimized by minimizing the reconstruction errors using the standard gradient descent. Specifically, the loss $\mathcal{L}$ is given by
$\mathcal{L}=(1-\rho)\mathcal{L}_{1}+\rho\mathcal{L}_{\text{D-SSIM}},$ (4)
where $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{D-SSIM}}$ are the $\ell_{1}$ norm and the structural dissimilarity between the rendered output and the ground truth, respectively. $\rho$ is a parameter that controls the tradeoff between the two losses.
3.2 Problem Formulation: Simplification as Constrained Optimization
To mitigate irreversible information loss caused by existing one-shot criterion-based pruning, our key innovation is to gradually impose substantial sparsity onto the Gaussians and maximally preserve information in the training process. Mathematically, considering a general 3DGS training, we introduce a sparsity constraint on the Gaussians to the training objective, $\min\mathcal{L}$. Thus, the 3DGS training is reformulated as another constrained optimization problem, the primary formulation of our proposed simplification framework, i.e.,
$\min{}$ $\mathcal{L},$ (5)
s.t. $\mathcal{N}(\bm{G})\leq\kappa.$
The introduced constraint $\mathcal{N}(\cdot)\leq\kappa$ restricts the number of Gaussians no greater than a target number $\kappa$.
However, the above problem cannot be solved because the Gaussians are continuous functions instead of differentiable variables. Recalling the rendering process with Eq. 3, the contribution of the $i$-th Gaussians to a particular pixel is finally determined by its opacity $a_{i}$. Alternatively, the original sparsity constraint directly on the Gaussians can be explicitly transformed into a sparsity constraint on the opacity variables. As each Gaussian only has one opacity variable, we can use a vector $\bm{a}$ to represent all the opacities and easily restrict its number of non-zeros with $\ell_{0}$ norm, $\|\bm{a}\|_{0}$. For better presentation, we use a variable set $\bm{\Theta}$ to represent other GS variables except for the opacity $\bm{a}$. Hence, the constrained optimization problem can be reformulated as
$\min_{\bm{a},\bm{\Theta}}{}$ $\mathcal{L}(\bm{a},\bm{\Theta}),$ (6)
s.t. $\|\bm{a}\|_{0}\leq\kappa.$
With this reformulation, we have simplified the primary problem to another form that our proposed “optimizing-sparsifying" can efficiently solve.
3.3 “Optimizing-Sparsifying” Solution
Figure 4: Evolution of opacity distribution and PSNR for 3DGS and GaussianSpa. The “optimizing-sparsifying" starts at iteration 15K. At iteration 25K, GaussianSpa removes “zero" Gaussians, followed by a light tuning.
Figure 5: Visual quality results on the Drjohnson and Counter scenes, compared to existing simplification approaches and vanilla 3DGS. The numbers of remaining Gaussians are displayed. It is observed that our GaussianSpa recovers details closest to the ground truth in the actual rendering outcomes with a significantly reduced number of Gaussians.
In this subsection, we present an efficient “optimizing-sparsifying" solution for problem Eq. 6. The key idea is to split it into two simple sub-problems separately solved in an “optimizing" step and a “sparsifying" step alternately.
In order to deal with the non-convex constraint, $\|\bm{a}\|_{0}\leq\kappa$, we first introduce an indicator function,
$h(\bm{a})=\left\{\begin{matrix}0,&\|\bm{a}\|_{0}\leq\kappa,\\ +\infty,&\textup{otherwise},\\ \end{matrix}\right.$ (7)
to the minimization objective and remove the sparsity constraint. Thus, problem Eq. 6 becomes an unconstrained optimization form, i.e.,
$\min_{\bm{a},\bm{\Theta}}~{}\mathcal{L}(\bm{a},\bm{\Theta})+h(\bm{a}).$ (8)
Here, the first term is the original 3DGS training objective, which is easy to handle, while the second term is still non-differentiable. This makes the above problem unable to solve using gradient descent under existing training frameworks such as PyTorch [44] and TensorFlow [1]. To decouple the two terms, we further introduce an auxiliary variable $\bm{z}$, whose size is identical to $\bm{a}$, and convert Eq. 8 as
$\min_{\bm{a},\bm{z},\bm{\Theta}}{}$ $\mathcal{L}(\bm{a},\bm{\Theta})+h(\bm{z}),$ (9)
s.t. $\bm{a}=\bm{z}.$
The above problem becomes another constrained problem with an equality constraint. Thus, the corresponding dual augmented Lagrangian [7] form can be given by
$L(\bm{a},\bm{z},\bm{\Theta},\bm{\lambda};\delta)=\mathcal{L}(\bm{a},\bm{\Theta})+h(\bm{z})+\frac{\delta}{2}\|\bm{a}-\bm{z}+\bm{\lambda}\|^{2}+\frac{\delta}{2}\|\bm{\lambda}\|^{2},$ (10)
where $\bm{\lambda}$ is the dual Lagrangian multiplier, and $\delta$ is the penalty parameter. Correspondingly, we can separately optimize over the 3DGS variables, $\bm{a},\bm{\Theta}$, and the auxiliary variable, $\bm{z}$, in the augmented Lagrangian, then solve two split sub-problems in an “optimizing" step and a “sparsifying" step alternately.
[…]
In the “optimizing" step, the minimization objective, Eq. 11, targets only 3DGS variables and contains two terms, which are the original 3DGS loss function and a quadratic regularization that enforces the opacity $\bm{a}$ close to the exactly sparse variable $\bm{z}$. Since both terms are differentiable, the sub-problem in this step can be directly optimized with gradient descent. In other words, the solution for this sub-problem is consistent with optimizing 3DGS variables with additional gradients, which are computed by
$\frac{\partial L}{\partial\bm{a}}$ $=\frac{\partial\mathcal{L}(\bm{a},\bm{\Theta})}{\partial\bm{a}}+\delta(\bm{a}-\bm{z}+\bm{\lambda}),$ (12)
$\frac{\partial L}{\partial\bm{\Theta}}$ $=\frac{\partial\mathcal{L}(\bm{a},\bm{\Theta})}{\partial\bm{\Theta}}.$ (13)
Hence, the opacity $\bm{a}$ and other GS variables $\bm{\Theta}$, can be updated by
$\bm{a}\leftarrow\bm{a}-\eta\frac{\partial L}{\partial\bm{a}},~{}~{}\bm{\Theta}\leftarrow\bm{\Theta}-\eta\frac{\partial L}{\partial\bm{\Theta}},$ (14)
respectively. Here, $\eta$ is the learning rate defined in the training process.
Remark-1. The “optimizing" step essentially optimizes over the original Gaussians and pushes the opacity $\bm{a}$ close to the auxiliary $\bm{z}$ that is exactly sparse using a gradient-based approach. This way, it can simultaneously satisfy the 3DGS performance and the sparsity requirements, achieving a sweet point in the training process.
[…]
In the “sparsifying" step, the minimization objective, Eq. 15, targets only the auxiliary variable, $\bm{z}$, while the term of indicator function dominates Eq. 15. This sub-problem essentially has a form of the proximal operator [43], $\mathrm{\bf{prox}}_{h}$, associated with the indicator function $h(\cdot)$. Thus, the solution of this sub-problem can be directly given by
$\bm{z}\leftarrow\mathrm{\bf{prox}}_{h}(\bm{a}+\bm{\lambda}).$ (16)
The proximal operator maps the opacity $\bm{a}$ to the auxiliary variable $\bm{z}$, which contains at most $\kappa$ non-zeroes. This operator results in a projection solution [8], sparsifying $(\bm{a}+\bm{\lambda})$, by projecting it onto a set that satisfies $\|\bm{z}\|_{0}\leq\kappa$. More specifically, the projection keeps top-$\kappa$ elements and sets the rest to zeros.
Remark-2. The “sparsifying" step essentially prunes the auxiliary variable $\bm{z}$, which can be considered the exactly sparse version of $\bm{a}$. By not directly operating the original Gaussians, this step avoids the irreversible loss of important information.
[…]
$\bm{\lambda}\leftarrow\bm{\lambda}+\bm{a}-\bm{z}.$ (17)
Overall, the optimization steps and the multiplier update are performed alternately in the 3DGS training process until convergence, as summarized in Algorithm 1. Generally, we consider our solution convergences once the condition, $\|\bm{a}-\bm{z}\|^{2}\leq\epsilon$, is satisfied, or the iterations reach the predefined maximum number.
Remark-3. In summary, our “optimizing-sparsifying" solution does not explicitly prune the Gaussians but temporarily “marks" them to be pruned on the introduced auxiliary variable $\bm{z}$. As those steps are performed alternately, summarized in Algorithm 1, the process eventually converges to a sparse set that maximizes the performance. Upon the convergence, the information can be preserved and transferred to the “non-zero" Gaussians. Another benefit over existing heuristic one-shot pruning approaches is that our removal of Gaussians does not rely on a one-time importance measurement but a long-term training process. As shown in Figure 2, existing one-shot, criterion-based methods lead to a sudden performance drop, followed by irreversible losses of 0.2 dB and 0.5 dB for importance-based and opacity-based methods, respectively, even with additional training iterations. In contrast, our GaussianSpa approach smoothly and optimally preserves knowledge, enabling more accurate view rendering. Moreover, the introduced variables, $\bm{z}$ and $\bm{\lambda}$, are considered constants, as we do not calculate their gradients. The “sparsifying" step is performed every tens iterations instead of every single iteration. Thus, the overall training speed is very similar to the original 3DGS.
Algorithm 1 Procedure of “Optimizing-Sparsifying"
Input: Gaussian opacity $\bm{a}$, 3DGS variables $\bm{\Theta}$, target number of Gaussians $\kappa$, penalty parameter $\delta$, feasibility tolerance $\epsilon$, maximum iterations $T$.
Output: Optimized $\bm{a}$ and $\bm{\Theta}$.
[…]
2:$t$leftarrow00;
3:while $\|\bm{a}-\bm{z}\|^{2}>\epsilon$ and $t\leq T$ do
4:     Update $\bm{a}$ and $\bm{\Theta}$ via Eq. 14; $\triangleright$ “Optimizing" Step
5:     Update $\bm{z}$ via Eq. 16; $\triangleright$ “Sparsifying" Step
6:     Update $\bm{\lambda}$ via Eq. 17; $\triangleright$ Multiplier Update
[…]
8:end while
Method Mip-NeRF 360 Tanks&Temples Deep Blending
PSNR↑ SSIM↑ LPIPS↓ #G/M↓ PSNR↑ SSIM↑ LPIPS↓ #G/M↓ PSNR↑ SSIM↑ LPIPS↓ #G/M↓
3DGS [25] 27.45 0.810 0.220 3.110 23.63 0.850 0.180 1.830 29.42 0.900 0.250 2.780
CompactGaussian [29] 27.08 0.798 0.247 1.388 23.32 0.831 0.201 0.836 29.79 0.901 0.258 1.060
LP-3DGS-R [54] 27.47 0.812 0.227 1.959 23.60 0.842 0.188 1.244 - - - -
LP-3DGS-M [54] 27.12 0.805 0.239 1.866 23.41 0.834 0.198 1.116 - - - -
EAGLES [22] 27.23 0.809 0.238 1.330 23.37 0.840 0.200 0.650 29.86 0.910 0.250 1.190
Mini-Splatting [20] 27.40 0.821 0.219 0.559 23.45 0.841 0.186 0.319 30.05 0.909 0.254 0.397
Taming 3DGS [34] 27.31 0.801 0.252 0.630 23.95 0.837 0.201 0.290 29.82 0.904 0.260 0.270
CompGS [38] 27.12 0.806 0.240 0.845 23.44 0.838 0.198 0.520 29.90 0.907 0.251 0.550
30.37 0.914 0.249 0.335
GaussianSpa 27.85 0.825 0.214 0.547 23.98 0.852 0.180 0.269 30.33 0.912 0.254 0.256
Table 1: Quantitative results on multiple datasets, compared with existing state-of-the-art works. Mini-Splatting [20] results are replicated using official code. 3DGS results are reported from [22]. “#G/M” represents number of million Gaussians.
3.4 Overall Workflow of GaussianSpa
With the proposed “optimizing-sparsifying"-integrated training process presented in the previous subsection, the Gaussians exhibit substantial sparsity characteristics. In other words, a certain number of Gaussian opacities are close to zeros, meaning those Gaussians have almost no contribution to the rendering, which can be directly removed. As the redundant Gaussians are appropriately removed, our simplified 3DGS model can exhibit superior performance after a light tuning, even slightly higher than the original 3DGS. The overall workflow of our GaussianSpa is illustrated in Figure 3.
We have visualized the evolution of Gaussian opacity distribution along with the PSNR, trained on the Room scene, as shown in Figure 4. It is observed that the opacity distribution of GaussianSpa is significantly distinct from the original 3DGS – there is a clear gap between the “zero" Gaussians and the remaining Gaussians in GaussianSpa, while the PSNR keeps increasing. This means GaussianSpa has successfully imposed a substantial sparsity property and preserved information to the “non-zero" Gaussians. At iteration 25K, we remove all “zero" Gaussians and perform a light tuning that further improves the performance, finally obtaining a compact model with high-quality rendering.
4 Experiments
4.1 Experimental Settings
Datasets, Baselines, and Metrics. We evaluate the rendering performance on multiple real-world datasets, including Mip-NeRF 360 [5], Tanks&Temples [27], and Deep Blending [23]. On all datasets, we compare our proposed GaussianSpa with vanilla 3DGS [25] and existing state-of-the-art Gaussian simplification approaches, including CompactGaussian [29], LP-3DGS [54], EAGLES [22], Mini-Spatting [20], Taming 3DGS [34], and CompGS [38]. We follow the standard practice to quantitatively evaluate the rendering performance, reporting results using the peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and learned perceptual image patch similarity (LPIPS). In addition to quantitative results, we comprehensively evaluate GaussianSpa with visual analyses such as ellipsoid view distribution and point cloud visualization against the existing state-of-the-art approaches.
Implementation Details. We conduct experiments under the same environment specified in the original 3DGS [25] with the PyTorch framework. Our experimental server has two AMD EPYC 9254 CPUs and eight NVIDIA GTX 6000 Ada GPUs. We use the same checkpoint files as Mini-Splatting [20] for training, which are stored before our “optimizing-sparsifying” process to ensure fair comparisons. GaussianSpa starts “optimizing-sparsifying” at iteration 15K and removes “zero" Gaussians at iteration 25K.
4.2 Quantitative Results
The quantitative results are summarized in Table 1, compared with various existing approaches. The table shows that our proposed GaussianSpa outperforms the baseline 3DGS across all three metrics with substantially fewer Gaussians. To be specific, after reducing the Gaussian points by 6$\times$, GaussianSpa still improves the PSNR by 0.4 dB over the vanilla 3DGS. Compared to other criterion-based simplification methods, including EAGLES [22], Mini-Splatting [20], Taming 3DGS [34], and CompGS [38], as well as the learnable mask-based approaches, CompactGaussia [29] and LP-3DGS [54], our GaussianSpa exhibits a significant improvement under a less number of Gaussians. Particularly, GaussianSpa achieves a PSNR improvement as high as 0.7 dB with even 3$\times$ fewer Gaussians, compared against LP-3DGS [54]. These results quantitatively demonstrate the superiority of our GaussianSpa.
We also plot quality curves with respect to the Gaussian reduction rate on the Room and Kitchen scenes in Figure 7 (more results are in Figure 13 in the Appendix), showing the robustness of GaussianSpa for different numbers of remaining Gaussians. Figure 7 shows that GaussianSpa outperforms the state-of-the-art method, Mini-Splatting [20], by an average of 0.5 dB across multiple Gaussian reduction rates.
Figure 6: Visualization for rendered Gaussian ellipsoids. GaussianSpa adaptively reconstructs high-frequency (e.g., air outlet of the train, complex textures of the carpet) and low-frequency (e.g., large-area sky, simple textures of the carpet) regions using appropriate numbers of Gaussians, respectively, more precisely matching the original textures and shapes, compared to the state-of-the-art Mini-Splatting [20] and the vanilla 3DGS.
4.3 Visual Quality Results
Figure 7: Quality-Reduction Rate curves on the (left) Kitchen and (right) Room scenes. GaussianSpa outperforms Mini-Splatting [20] across all reduction rates.
As depicted in Figure 5, we present the rendered images on the Drjohnson and Counter scenes [5, 23] for visual quality evaluation, in comparison to LightGaussian [19], Mini-Splattin [20], and the vanilla 3DGS. Together with Figure 1, those outcomes demonstrate the superior rendering quality of our GaussianSpa on diverse scenes over existing approaches, consistent with the quantitative results in the previous subsection. For instance, as shown in Figure 5, GaussianSpa accurately renders the outlet behind the railings for the Playroom scene, while LightGaussian [19] and Mini-Splatting [20] lose this object. Furthermore, GaussianSpa fully recovers the original details of the zoom-in object in the Counter scene, while other methods outcome blurry surfaces.
We comprehensively analyze the proposed GaussianSpa by visualizing the rendered Gaussian ellipsoids and final point clouds in Figure 6 and Figure 8, respectively. Those visualized results demonstrate that GaussianSpa can adaptively control the density to match the details when sparsifying the Gaussians, ensuring high-quality sparse 3D representation. From Figure 6, we can observe that GaussianSpa uses fewer but larger Gaussians to represent the blue sky and uses dense Gaussians to render complex carpet textures, providing precise rendering of details in the zoom-in images. In contrast, Mini-Splatting [20] cannot fully reconstruct the details, leading to blurry rendered textures. Figure 8 further demonstrates our superiority in simplification among existing approaches. GaussianSpa adaptively reduce the Gaussians in the low-frequency areas and preserve more Gaussians in high-frequency areas that generally need more Gaussians to synthesize, resulting in sparse Gaussian point clouds that precisely outline the scene contours.
5 Discussion
5.1 Convergence Analysis
Figure 8: Visualized point clouds for the Room and Drjohnson scenes. The numbers of remaining Gaussians in millions are displayed. With GaussianSpa, sparse Gaussians concentrate on high-frequency areas after our “optimizing-sparsifying"-integrated training process, ensuring the capability to capture detail-rich textures and shapes. On the contrary, the point clouds produced by other approaches cannot outline the contours.
We analyze convergence behavior of the proposed GaussianSpa and appropriately set hyper-parameters, such as $\delta$, which controls the sparsity strength in Eq. 10. As introduced in Section 3, GaussianSpa performs the “sparsifying" step every fixed number of iterations, referred to here as “interval" here. We plot loss curves for multiple $\delta$’s and intervals in Figure 9. It is seen that with different $\delta$ and interval values, the curves converge well at a similar rate during our “optimizing-sparsifying"-integrated training process. After removing “zero" Gaussians at iteration 25K, the loss curve still exhibits a consistent convergence, confirming the feasibility of GaussianSpa.
Figure 9: Loss curves with multiple (left) $\delta$ and (right) interval settings. The curves show that GaussianSpa exhibits a good convergence behaviour in the “optimizing-sparsifying"-integrated training process.
5.2 Generality of GaussianSpa
GaussianSpa is a general Gaussian simplification framework. Existing importance measurement criteria can be applied to project the auxiliary variable $\bm{z}$ onto the sparse space in the “sparsifying" step, Eq. 15, obtaining more compact 3DGS models with improved rendering quality. For example, we employ the criterion from LightGaussian [19] and plot the PSNR curves at multiple Gaussian reduction rates in Figure 13 in the Appendix. It is observed that GaussianSpa consistently outperforms LightGaussian [19] with an average of 0.4 dB improvement for the Playroom and Kitchen scenes.
5.3 Storage Compression
While GaussianSpa focuses on simplifying and optimizing the number of Gaussians to obtain a sparse 3DGS representation, existing compression techniques such as SH optimization and vector quantization can be applied as add-ons to the simplified 3DGS model. As shown in Table 2, we present the final storage cost for the Mip-NeRF 360 dataset by incorporating the SH distillation and vector quantization methods proposed by LightGaussian [19]. This comparison demonstrates that GaussianSpa can achieve the lowest storage consumption with additional compression techniques while maintaining superior rendering quality over existing hybrid compression methods, such as EfficientGS [31] and LightGaussian [19].
5.4 Relation to Other Optimization Algorithms
The “optimizing-sparsifying" algorithm in our proposed GaussianSpa can be categorized into the class of optimization algorithms that project the target variable onto a constrained space, specifically the $\ell_{p}$ ball (where $p=0$ in our case). This approach is similar to foundational optimization algorithms in compressed sensing [16] and sparse learning [9], such as Ridge regression [24] and Lasso regression [48]. These algorithms extend the least squares regression by incorporating $\ell_{p}$ regularization – Ridge regression adds the $\ell_{2}$ regularization, while the Lasso regression adds the $\ell_{1}$ regularization. The regularization acts as a penalty, combined with quadratic loss, reformulating the problem as a convex optimization task.
Method PSNR↑ SSIM↑ LPIPS↓ Storage↓
EfficientGS [31] 27.38 0.817 0.216 98 MB
LightGaussian [19] 27.28 0.805 0.243 42 MB
GaussianSpa 27.85 0.825 0.214 25 MB
Table 2: Storage comparison evaluated on the Mip-NeRF 360 dataset. GaussianSpa’s storage cost is reported based on the add-on compression methods (i.e., SH distillation and vector quantization) from LightGaussian [19].
However, neither $\ell_{1}$ nor $\ell_{2}$ regularization yields truly sparse solutions; achieving genuine sparsity requires the $\ell_{0}$ regularization. Yet, when $p=0$, the problem, also known as the sparse coding problem [41], becomes nonconvex and is generally recognized as NP-hard [15, 37, 49]. This complexity makes it computationally impractical when the dimension of the target variable is high. In such cases, Lasso (where $p=1$) offers a convex constraint region, serving as the tightest convex surrogate for the sparse coding problem, albeit with some sacrifice in exact sparsity [11]. Another alternative solver suitable for this case is a greedy algorithm, such as the Matching Pursuit algorithm [33, 45]. However, this algorithm cannot guarantee the global optima.
Iterative thresholding (IT) is another direction to address this problem effectively [17, 14, 18]. These algorithms perform a matrix multiplication and its adjoint at each iteration, followed by a scalar shrinkage step, which yields a simple yet highly effective structure. Furthermore, a closed-form solution for the global minimizer of the sparse coding problem can be obtained – even for non-convex cases – under the condition that a unitary matrix is involved [17]. Inspired by this, we introduce a duplicated variable (i.e., $\bm{z}$) and project it onto the $\ell_{0}$ ball, optimizing over two variables alternately, similar to the alternating direction method of multipliers [8]. By leveraging the simplicity of the unitary case, the “sparsifying" step aligns with this scenario and thus admits an explicit solution. Hence, our algorithm simultaneously benefits from the sparsity imposed by the $\ell_{0}$ constraint and achieves fast convergence, guided by the explicit solution.
6 Conclusion
In this paper, we present an optimization-based 3DGS simplification framework, GaussianSpa, for compact and high-quality sparse view synthesis. GaussianSpa formulates 3DGS simplification as a constrained optimization problem with a sparsity constraint on the Gaussian opacities. Then, we propose an “optimizing-sparsifying" solution to efficiently solve the formulated problem during the 3DGS training process. We comprehensively evaluate the proposed GaussianSpa on multiple datasets with quantitative results and qualitative analyses, demonstrating the superiority in rendering quality with fewer Gaussians compared to existing state-of-the-art approaches.
Appendix
7 Additional Results
7.1 Additional Quantitative Results
We summarize additional quantitative results on the Mip-NeRF 360, Tanks&Temples, and Deep Blending datasets in Table 4, Table 5, and Table 3, respectively. We plot additional PSNR-#Gaussians curves on diverse Scenes in Figure 13, in comparison with Mini-Splatting [20] and LightGaussian [19]. It can be observed that with the same number of Gaussians, our GaussianSpa shows superior rendering outcomes.
7.2 Additional Visual Quality Results
Figure 10 and Figure 11 provide additional rendered images comparing GaussianSpa with LightGaussian [19], Mini-Splatting [20], and original 3DGS on various scenes. Those additional results demonstrate our GaussianSpa achieves stronger representational power for background and detail-rich areas such as walls, carpets, and ladders, showcasing superior rendering qualities. We also present additional point cloud views in Figure 12. This further illustrates that GaussianSpa creates a high-quality sparse 3D representation that adaptively uses more Gaussians to represent high-frequency areas.
Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
Drjohnson 3DGS 28.77 0.900 0.250 3.260
Mini-Splatting 29.37 0.904 0.261 0.377
GaussianSpa 29.89 0.913 0.243 0.450
GaussianSpa 29.82 0.909 0.254 0.293
Playroom 3DGS 30.07 0.900 0.250 2.290
Mini-Splatting 30.72 0.914 0.248 0.417
GaussianSpa 30.84 0.916 0.254 0.219
GaussianSpa 30.84 0.916 0.254 0.219
Average 3DGS 29.42 0.900 0.250 2.780
Mini-Splatting 30.05 0.909 0.254 0.397
GaussianSpa 30.37 0.914 0.249 0.335
GaussianSpa 30.33 0.912 0.254 0.256
Table 3: Deep Blending per scene results. 3DGS results are reported from [22]. Mini-Splatting [20] results are replicated using official code.
Figure 10: Additional visual comparison on more scenes. The numbers of remaining Gaussians in millions are displayed.
Figure 11: (Continue) Additional visual comparison on more scenes. The numbers of remaining Gaussians in millions are displayed.
Figure 12: Additional visualized point clouds. The numbers of remaining Gaussians in millions are displayed.
Figure 13: The first 13 sub-figures: Quality-#G (the number of Gaussians in millions) curves comparing GaussianSpa with Mini-Splatting [20] on multiple scenes. Our GaussianSpa consistently outperforms Mini-Splatting [20] with the same #G. The last two sub-figures: Quality-Reduction Rate curves comparing GaussianSpa with LightGaussian [19] on the Kitchen and Playroom scenes.
Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
Bicycle 3DGS 25.13 0.750 0.240 5.310
Mini-Splatting 25.21 0.760 0.247 0.646
GaussianSpa 25.44 0.769 0.246 0.656
Bonsai 3DGS 32.19 0.950 0.180 1.250
Mini-Splatting 31.73 0.945 0.180 0.360
GaussianSpa 32.40 0.947 0.174 0.372
Counter 3DGS 29.11 0.910 0.180 1.170
Mini-Splatting 28.53 0.911 0.184 0.408
GaussianSpa 29.23 0.919 0.176 0.392
Flowers 3DGS 21.37 0.590 0.360 3.470
Mini-Splatting 21.42 0.616 0.336 0.670
GaussianSpa 21.75 0.610 0.329 0.674
Garden 3DGS 27.32 0.860 0.120 5.690
Mini-Splatting 26.99 0.842 0.156 0.738
GaussianSpa 27.26 0.848 0.151 0.728
Kitchen 3DGS 31.53 0.930 0.120 1.770
Mini-Splatting 31.24 0.929 0.122 0.438
GaussianSpa 32.03 0.934 0.117 0.423
Room 3DGS 31.59 0.920 0.200 1.500
Mini-Splatting 31.44 0.929 0.193 0.394
GaussianSpa 32.04 0.933 0.188 0.355
Stump 3DGS 26.73 0.770 0.240 4.420
Mini-Splatting 27.35 0.803 0.219 0.717
GaussianSpa 27.56 0.808 0.218 0.690
Treehill 3DGS 22.61 0.640 0.350 3.420
Mini-Splatting 22.69 0.652 0.332 0.663
GaussianSpa 22.94 0.660 0.329 0.637
Average 3DGS 27.45 0.810 0.220 3.110
Mini-Splatting 27.40 0.821 0.219 0.559
GaussianSpa 27.85 0.825 0.214 0.547
Table 4: MiP-NeRF360 per scene results. 3DGS results are reported from [22]. Mini-Splatting [20] results are replicated using official code.
Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
Train 3DGS 21.94 0.810 0.200 1.110
Mini-Splatting 21.78 0.805 0.231 0.287
GaussianSpa 22.17 0.815 0.228 0.199
Truck 3DGS 25.31 0.880 0.150 2.540
Mini-Splatting 25.13 0.878 0.141 0.352
GaussianSpa 25.79 0.888 0.132 0.338
Average 3DGS 23.63 0.850 0.180 1.830
Mini-Splatting 23.45 0.841 0.186 0.319
GaussianSpa 23.98 0.852 0.180 0.269
Table 5: Tanks&Temples per scene results. 3DGS results are reported from [22]. Mini-Splatting [20] results are replicated using official code.

### v3

@@
-Yangming Zhang*    Wenqi Jia* Dept. of Computer Science
+GaussianSpa: An “Optimizing-Sparsifying" Simplification Framework for Compact and High-Quality 3D Gaussian Splatting
 Abstract
@@
 3D Gaussian Splatting (3DGS) has emerged as a mainstream for novel view synthesis, leveraging continuous aggregations of Gaussian functions to model scene geometry.
-However, 3DGS suffers from substantial memory requirements to store the multitude of Gaussians, hindering its practicality.
+However, 3DGS suffers from substantial memory requirements to store the large amount of Gaussians, hindering its efficiency and practicality.
 To address this challenge, we introduce GaussianSpa, an optimization-based simplification framework for compact and high-quality 3DGS.
-Specifically, we formulate the simplification as an optimization problem associated with the 3DGS training.
-Correspondingly, we propose an efficient “optimizing-sparsifying” solution that alternately solves two independent sub-problems, gradually imposing strong sparsity onto the Gaussians in the training process.
-Our comprehensive evaluations on various datasets show the superiority of GaussianSpa over existing state-of-the-art approaches.
+Specifically, we formulate the simplification objective as a constrained optimization problem associated with the 3DGS training.
+Correspondingly, we propose an efficient “optimizing-sparsifying” solution for the formulated problem, alternately solving two independent sub-problems and gradually imposing substantial sparsity onto the Gaussians in the 3DGS training process.
+We conduct quantitative and qualitative evaluations on various datasets, demonstrating the superiority of GaussianSpa over existing state-of-the-art approaches.
 Notably, GaussianSpa achieves an average PSNR improvement of 0.9 dB on the real-world Deep Blending dataset with 10$\times$ fewer Gaussians compared to the vanilla 3DGS.
@@
 Figure 1: We present GaussianSpa, enabling high-quality and compact view synthesis with superior rendering of details.
-Compared to the existing state-of-the-art method, Mini-Splatting [20], our GaussianSpa captures detail-rich textures and background more accurately with fewer Gaussians.
+Compared to the existing state-of-the-art method, Mini-Splatting [17], our GaussianSpa captures detail-rich textures and low-frequency background more accurately with fewer Gaussians.
 1 Introduction
-Novel view synthesis has become a pivotal area in computer vision and graphics, driving advancements in applications such as virtual reality, augmented reality, and immersive media experiences [21].
+Novel view synthesis has become a pivotal area in computer vision and graphics, advancing applications such as virtual reality, augmented reality, and immersive media experiences [18].
 NeRF [32] has recently gained prominence in this domain because it can generate high-quality, photorealistic images from sparse input views by representing scenes as continuous volumetric functions based on neural networks.
-However, NeRF often requires extensive computational resources and long training times, making it less practical for real-time applications and large-scale reconstructions.
+However, NeRF requires substantial computational resources and long training times, making it less practical for real-time and large-scale applications.
 3D Gaussian Splatting (3DGS) [22] has emerged as a powerful alternative, leveraging continuous aggregations of Gaussian functions to model scene geometry and appearance.
 Unlike NeRF, which relies on neural networks to approximate volumetric radiance fields, 3DGS directly represents scenes using a collection of Gaussians.
-This approach excels in capturing details and smooth transitions, offering significant advantages in training and rendering speed.
+This approach effectively captures details and smooth transitions, offering faster training and rendering.
 3DGS achieves superior visual fidelity [25] compared to NeRF while reducing computational overhead, making it more suitable for interactive applications that demand quality and performance.
 Despite its strengths, 3DGS suffers from significant memory requirements that hinder its practicality.
-The main issue is the massive memory consumption associated with storing a large number of Gaussians that are needed to represent complex scenes.
-Each Gaussian occupies memory space for its parameters, including position, covariance, and color attributes.
+The main issue is the massive memory consumption required to store numerous Gaussians, each with parameters like position, covariance, and color.
 In densely sampled scenes, the sheer volume of Gaussians leads to memory usage that exceeds the capacity of typical hardware, making it challenging to handle higher-resolution scenes and limiting its applicability in resource-constrained environments.
@@
 Techniques such as pruning and sampling aim to discard unimportant Gaussians based on hand-crafted criteria such as opacity [22, 50], importance score (hit count) [17, 16, 37], dominant primitives [28], and binary mask [51, 26].
-However, these criteria are generally utilized to determine the importance of Gaussian points from a single heuristic perspective, limiting robustness in dynamic scenes or under varying lighting conditions.
-Moreover, the sudden one-shot removal may cause permanent loss of Gaussians that are crucial to visual synthesis, making it challenging to recover the original performance after even long-term training, as shown in Figure 2.
+However, such single-perspective heuristic criteria may lack robustness in dynamic scenes or varying lighting.
+Moreover, the sudden one-shot removal may cause permanent loss of Gaussians crucial to visual synthesis, making it challenging to recover the original performance after even long-term training, as shown in Figure 2.
+Mask-based pruning methods also suffer from this issue due to weak sparsity enforcement on the multitude of Gaussians.
 Consequently, while these methods can alleviate memory and storage burdens to some extent, they often lead to sub-optimal rendering outcomes with loss of details and visual artifacts, thereby compromising the quality of the synthesized views.
-In this paper, we present an optimization-based simplification framework, GaussianSpa, for compact and high-quality Gaussian Splatting.
+Figure 2: PSNR curves of hand-crafted criteria-based pruning methods. Gaussians are removed by 85% at iteration 25K.
+In this paper, we present an optimization-based simplification framework, GaussianSpa, for compact and high-quality 3DGS.
 In the proposed framework, we formulate 3DGS simplification as a constrained optimization problem under a target number of Gaussians.
@@
 In the “sparsifying" step, we analytically project the auxiliary Gaussians onto the constrained sparse space.
-  • We comprehensively evaluate GaussianSpa through extensive experiments on various complex scenes, demonstrating improved rendering quality compared to existing approaches.
-Particularly, with as high as 10$\times$ fewer number of Gaussians than the vanilla 3DGS, GaussianSpa achieves an average 0.4 dB improvement on the Mip-NeRF 360 [4] and Tanks&Temples [27] datasets, 0.9 dB on Deep Blending [23] dataset.
-Furthermore, we conduct various visual quality assessments, showing that GaussianSpa exhibits high-quality rendering of details and sparse 3D Gaussian views.
+  • We comprehensively evaluate GaussianSpa via extensive experiments on various complex scenes, demonstrating improved rendering quality compared to existing state-of-the-art approaches.
+Particularly, with as high as 10$\times$ fewer number of Gaussians than the vanilla 3DGS, GaussianSpa achieves an average 0.4 dB PSNR improvement on the Mip-NeRF 360 [2] and Tanks&Temples [24] datasets, 0.9 dB on Deep Blending [20] dataset.
+Furthermore, we conduct various visual quality analyses, demonstrating the superiority of GaussianSpa in rendering high-quality views and capturing rich details.
 2 Related Work
+Figure 3: Overall workflow of our proposed GaussianSpa framework.
 2.1 Novel View Synthesis
-Novel View Synthesis (NVS) [2] is a technique in computer vision and graphics to generate images of a 3D scene from unseen viewpoints based on existing multi-view data.
+Novel View Synthesis (NVS) [1] generates images of a 3D scene from unseen viewpoints based on existing multi-view data.
 A seminal work in this field is NeRF [32], which uses an implicit neural network to represent a scene as a continuous 5-dimensional function, mapping 3D coordinates and viewing directions to color and density.
-NeRF introduces positional encoding and a hierarchical sampling strategy to improve the ability of the MLP [6] to represent high-frequency image details and accelerate the rendering process.
-However, NeRF’s reliance on single narrow-ray sampling poses challenges in addressing aliasing and blur artifacts.
-To mitigate this, Mip-NeRF [3] represents the scene at a continuous-valued scale, enabling anti-aliased rendering by casting conical frustums from each pixel at various scales.
-Additionally, Mip360 [4] extends NeRF’s capabilities to unbounded scenes, utilizing a non-linear scene parameterization, online distillation, and a novel distortion-based regularizer.
-More recently, an explicit point-based method, 3D Gaussian Splatting (3DGS), which leverages continuous aggregations of Gaussian functions to model scene geometry, has emerged as a point-based alternative, demonstrating significant improvements in rendering quality and speed.
-Figure 2: PSNR curves of human-crafted criteria-based pruning methods. All methods remove 85% of the Gaussians at iteration 25K.
+However, its reliance on single narrow-ray sampling poses challenges in addressing aliasing and blur artifacts.
+More recently, 3D Gaussian Splatting (3DGS), which leverages continuous aggregations of Gaussian functions to model scene geometry, has emerged as a point-based alternative, demonstrating significant improvements in rendering quality and speed.
 2.2 Efficient 3DGS
@@
 However, the demands of real-time rendering and the reconstruction of unbounded, large-scale 3D scenes underscore the requirement for model efficiency [7].
-Optimization for 3DGS focuses on four main categories to effectively reduce model size: Gaussian simplification, vector quantization, spherical harmonics optimization, and hybrid compression.
+Optimization for efficient 3DGS focuses on four main categories: Gaussian simplification, quantization, spherical harmonics optimization, and hybrid compression.
 Gaussian Simplification.
@@
 Prior approaches [17, 37, 31, 51, 29, 43, 27, 10, 23] mainly focus on pruning redundant Gaussians according to hand-crafted importance criteria.
-Mini-Splatting [20] introduces blur splitting and depth reinitialization to address overlapping and under-reconstructed problems, using stochastic sampling to better preserve geometry based on importance scores.
-Radsplatting [40] computes an importance score by aggregating ray contributions of Gaussians across all input images, replacing summation with a max operator for robustness to varying scene coverage.
-Taming 3DGS [34] implements steerable densification by calculating scores based on pixel saliency, gradients, and Gaussian attributes, then selectively cloning or splitting high-scoring Gaussians.
-LP-3DGS [54] applies a trainable binary mask with Gumbel-Sigmoid to automatically optimize pruning, improving efficiency while maintaining quality.
-Vector Quantization.
-Vector quantization [50] aims to compress data using discrete entries from a learned codebook, reducing data redundancy.
-Several works [22, 38, 12, 51, 42, 36, 39] apply vector quantization for compressing 3DGS models.
-For instance, EAGLES [22] quantizes attributes like spherical harmonics (SH) coefficients and opacity to create latent representations for per-point compression, using an MLP decoder to reconstruct the original attributes with high fidelity.
-CompGS [38] applies a K-means-based vector quantization approach to compress 3DGS parameters, significantly reducing storage and rendering time with minimal quality loss.
-RDO-Gaussian [51] performs end-to-end rate-distortion optimization for 3D Gaussian Splatting, combining dynamic pruning and entropy-constrained vector quantization to reduce storage while maintaining quality.
+Mini-Splatting [17] addresses overlapping and reconstruction issues with blur splitting, depth reinitialization, and stochastic sampling.
+Radsplatting [37] enhances robustness by using a max operator to compute importance scores from ray contributions.
+Taming 3DGS [31] applies selective densification based on pixel saliency and gradients, while LP-3DGS [51] learns a binary mask for efficient pruning.
+Quantization.
+Quantization [47] aims to compress data using discrete entries without appreciable degradation in quality, reducing data redundancy.
+Several works [19, 35, 9, 48, 39, 33, 36] apply quantization for compressing 3DGS models.
+For example, EAGLES [19] quantizes attributes like spherical harmonics (SH) coefficients.
+Vector quantization, a form of quantization, compresses data by mapping vectors to a learned codebook.
+CompGS [35] leverages K-means-based vector quantization to reduce storage and rendering time, while RDO-Gaussian [48] combines pruning and entropy-constrained vector quantization for efficient compression with minimal quality loss.
 Spherical Harmonics Optimization.
-Among 3DGS parameters, spherical harmonics (SH) coefficients take up a significant portion, requiring (45+3) floating-point values per Gaussian for a degree of 3 (as used in [25]), which accounts for around 80% of the total attribute volume.
+Among 3DGS parameters, spherical harmonics (SH) coefficients take up a significant portion, requiring (45+3) floating-point values per Gaussian for a degree of 3 [22], which accounts for around 80% of the total attribute volume.
 Prior works [16, 26] optimize SH coefficients to reduce storage overhead.
@@
 In addition to methods focused on individual aspects of 3DGS compression, several recent approaches [16, 28, 26, 49] combine multiple techniques for hybrid compression to achieve further storage reduction.
-LightGaussian [19] prunes 3DGS points using a global importance score, applies vector quantization to compress SH coefficients, and reduces SH degree through knowledge distillation combined with pseudo-view augmentation.
-EfficientGS [31] not only keeps zeroth-order SH for all Gaussians, increasing the order only when needed but also classifies Gaussians into steady and non-steady states, densifying non-steady Gaussians while retaining only the top-K most important Gaussians per ray.
-CompactGaussian [29] introduces a learnable mask to mark redundant Gaussians and residual vector quantization for scale and rotation and replaces SH with a grid-based neural field to represent colors.
-Although these methods improve 3DGS efficiency, they lead to irreversible information loss during the Gaussian simplification process.
-Specifically, those methods perform simplification by a one-shot pruning on the Gaussians without prominent sparse characteristics, causing a notable PSNR degradation that is challenging to recover, as shown in Figure 2.
-Consequently, when applying additional compressions like SH optimization and vector quantization after simplification, the compressed 3DGS model often exhibits further performance loss due to the accumulated errors.
-Our work focuses on Gaussian simplification by formulating it as an optimization problem and imposing substantial sparsity onto Gaussians, mitigating performance issues with improved rendering quality at the first stage to achieve efficient 3DGS.
-Figure 3: Overall workflow of our proposed GaussianSpa framework.
+LightGaussian [16] uses a global importance score for pruning and applies vector quantization to compress SH coefficients.
+EfficientGS [28] selectively increases SH order and densifies non-steady Gaussians.
+CompactGaussian [26] introduces a learnable mask for pruning, uses residual vector quantization for scale and rotation, and replaces SH with a grid-based neural field for color representation.
+Although mask-based method impose sparsity relying on mask loss, they and other methods both lead to irreversible performance loss in the Gaussian simplification process by performing a one-shot pruning on the Gaussians, as shown in Figure 2.
+Consequently, performance often worsens when applying SH optimization and quantization due to accumulated errors.
+Our work focuses on Gaussian simplification by formulating it as an optimization problem and imposing substantial sparsity onto Gaussians, mitigating performance issues with improved rendering quality.
 3 Methodology: GaussianSpa
+Figure 4: Visual quality results on the Drjohnson and Counter scenes, compared to existing simplification approaches and vanilla 3DGS.
+The numbers of remaining Gaussians are displayed.
+It is observed that our GaussianSpa recovers details closest to the ground truth in the actual rendering outcomes with a significantly reduced number of Gaussians.
 3.1 Background of 3D Gaussian Splatting
@@
 The covariance matrix $\bm{\Sigma}$ is generally composed of $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$ with a rotation matrix $\bm{R}$ and a scale matrix $\bm{S}$, ensuring the positive definite property.
-Additionally, each Gaussian is associated with an opacity $a$ and spherical harmonics (SH) coefficients.
+Each Gaussian also has an associated opacity $a$ and spherical harmonics (SH) coefficients.
 When rendering a 2D image from the 3D scene at a certain camera angle, 3DGS projects 3D Gaussians to that 2D plane and calculates the projected 2D covariance matrix by
@@
 Here, $c_{i}$ and $\alpha_{i}$ denote the view-dependent color computed from the corresponding SH coefficients and the rendering opacity calculated by the Gaussian opacity $a_{i}$, respectively, corresponding to the $i$-th Gaussian.
-In the training process [25], the Gaussians are initialized from a sparse point cloud generated by Structure-from-Motion (SfM) techniques [47].
+In the training process [22], the Gaussians are initialized from a sparse point cloud generated by Structure-from-Motion (SfM) [44].
 Then, the number of Gaussians is adjusted by a densification algorithm [22] according to the reconstruction geometry.
-Afterward, each Gaussian’s variables, including the center position, rotation matrix, scaling matrix, opacity, and SH coefficients, are optimized by minimizing the reconstruction errors using the standard gradient descent.
+Afterward, each Gaussian’s attributes, including the center position, rotation matrix, scaling matrix, opacity, and SH coefficients, are optimized by minimizing the reconstruction errors using the standard gradient descent.
 Specifically, the loss $\mathcal{L}$ is given by
-$\mathcal{L}=(1-\rho)\mathcal{L}_{1}+\rho\mathcal{L}_{\text{D-SSIM}},$ (4)
+$\mathcal{L}=(1-\rho)\mathcal{L}_{1}+\rho\mathcal{L}_{\text{D-SSIM}},\vspace{-1mm}$ (4)
 where $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{D-SSIM}}$ are the $\ell_{1}$ norm and the structural dissimilarity between the rendered output and the ground truth, respectively.
@@
 The introduced constraint $\mathcal{N}(\cdot)\leq\kappa$ restricts the number of Gaussians no greater than a target number $\kappa$.
-However, the above problem cannot be solved because the Gaussians are continuous functions instead of differentiable variables.
+However, the above problem cannot be solved because the restriction on the number of Gaussian functions is not differentiable.
 Recalling the rendering process with Eq.
-3, the contribution of the $i$-th Gaussians to a particular pixel is finally determined by its opacity $a_{i}$.
+3, the contribution of the $i$-th Gaussian to a particular pixel is finally determined by its opacity $a_{i}$.
 Alternatively, the original sparsity constraint directly on the Gaussians can be explicitly transformed into a sparsity constraint on the opacity variables.
@@
 3.3 “Optimizing-Sparsifying” Solution
-Figure 4: Evolution of opacity distribution and PSNR for 3DGS and GaussianSpa.
-The “optimizing-sparsifying" starts at iteration 15K.
-At iteration 25K, GaussianSpa removes “zero" Gaussians, followed by a light tuning.
-Figure 5: Visual quality results on the Drjohnson and Counter scenes, compared to existing simplification approaches and vanilla 3DGS.
-The numbers of remaining Gaussians are displayed.
-It is observed that our GaussianSpa recovers details closest to the ground truth in the actual rendering outcomes with a significantly reduced number of Gaussians.
+Algorithm 1 Procedure of “Optimizing-Sparsifying"
+Input: Gaussian opacity $\bm{a}$, 3DGS variables $\bm{\Theta}$, target number of Gaussians $\kappa$, penalty parameter $\delta$, feasibility tolerance $\epsilon$, maximum iterations $T$.
+Output: Optimized $\bm{a}$ and $\bm{\Theta}$.
+[…]
+2:$t$leftarrow00;
+3:while $\|\bm{a}-\bm{z}\|^{2}>\epsilon$ and $t\leq T$ do
+4:     Update $\bm{a}$ and $\bm{\Theta}$ with Eq. 14; $\triangleright$ “Optimizing" Step
+5:     Update $\bm{z}$ with Eq. 16; $\triangleright$ “Sparsifying" Step
+6:     Update $\bm{\lambda}$ with Eq. 17; $\triangleright$ Multiplier Update
+[…]
+8:end while
 In this subsection, we present an efficient “optimizing-sparsifying" solution for problem Eq.
@@
 In order to deal with the non-convex constraint, $\|\bm{a}\|_{0}\leq\kappa$, we first introduce an indicator function,
-$h(\bm{a})=\left\{\begin{matrix}0,&\|\bm{a}\|_{0}\leq\kappa,\\ +\infty,&\textup{otherwise},\\ \end{matrix}\right.$ (7)
+$h(\bm{a})=\left\{\begin{matrix}0,&\|\bm{a}\|_{0}\leq\kappa,\\ +\infty,&\textup{otherwise},\\ \end{matrix}\right.\vspace{-1mm}$ (7)
 to the minimization objective and remove the sparsity constraint. Thus, problem Eq. 6 becomes an unconstrained optimization form, i.e.,
-$\min_{\bm{a},\bm{\Theta}}~{}\mathcal{L}(\bm{a},\bm{\Theta})+h(\bm{a}).$ (8)
+$\min_{\bm{a},\bm{\Theta}}~{}\mathcal{L}(\bm{a},\bm{\Theta})+h(\bm{a}).\vspace{-1mm}$ (8)
 Here, the first term is the original 3DGS training objective, which is easy to handle, while the second term is still non-differentiable.
-This makes the above problem unable to solve using gradient descent under existing training frameworks such as PyTorch [44] and TensorFlow [1].
+This prevents the above problem from being solved using gradient-based solutions under existing training frameworks such as PyTorch [41].
 To decouple the two terms, we further introduce an auxiliary variable $\bm{z}$, whose size is identical to $\bm{a}$, and convert Eq.
@@
 […]
+Figure 5: Evolution of opacity distribution and PSNR for 3DGS and GaussianSpa.
+The “optimizing-sparsifying" starts at iteration 15K.
+At iteration 25K, GaussianSpa removes “zero" Gaussians, followed by a light tuning.
 In the “optimizing" step, the minimization objective, Eq.
@@
 $\bm{a}\leftarrow\bm{a}-\eta\frac{\partial L}{\partial\bm{a}},~{}~{}\bm{\Theta}\leftarrow\bm{\Theta}-\eta\frac{\partial L}{\partial\bm{\Theta}},$ (14)
-respectively. Here, $\eta$ is the learning rate defined in the training process.
-Remark-1.
-The “optimizing" step essentially optimizes over the original Gaussians and pushes the opacity $\bm{a}$ close to the auxiliary $\bm{z}$ that is exactly sparse using a gradient-based approach.
-This way, it can simultaneously satisfy the 3DGS performance and the sparsity requirements, achieving a sweet point in the training process.
-[…]
-In the “sparsifying" step, the minimization objective, Eq.
-15, targets only the auxiliary variable, $\bm{z}$, while the term of indicator function dominates Eq.
-15.
-This sub-problem essentially has a form of the proximal operator [43], $\mathrm{\bf{prox}}_{h}$, associated with the indicator function $h(\cdot)$.
-Thus, the solution of this sub-problem can be directly given by
-$\bm{z}\leftarrow\mathrm{\bf{prox}}_{h}(\bm{a}+\bm{\lambda}).$ (16)
-The proximal operator maps the opacity $\bm{a}$ to the auxiliary variable $\bm{z}$, which contains at most $\kappa$ non-zeroes.
-This operator results in a projection solution [8], sparsifying $(\bm{a}+\bm{\lambda})$, by projecting it onto a set that satisfies $\|\bm{z}\|_{0}\leq\kappa$.
-More specifically, the projection keeps top-$\kappa$ elements and sets the rest to zeros.
-Remark-2.
-The “sparsifying" step essentially prunes the auxiliary variable $\bm{z}$, which can be considered the exactly sparse version of $\bm{a}$.
-By not directly operating the original Gaussians, this step avoids the irreversible loss of important information.
-[…]
-$\bm{\lambda}\leftarrow\bm{\lambda}+\bm{a}-\bm{z}.$ (17)
-Overall, the optimization steps and the multiplier update are performed alternately in the 3DGS training process until convergence, as summarized in Algorithm 1.
-Generally, we consider our solution convergences once the condition, $\|\bm{a}-\bm{z}\|^{2}\leq\epsilon$, is satisfied, or the iterations reach the predefined maximum number.
-Remark-3.
-In summary, our “optimizing-sparsifying" solution does not explicitly prune the Gaussians but temporarily “marks" them to be pruned on the introduced auxiliary variable $\bm{z}$.
-As those steps are performed alternately, summarized in Algorithm 1, the process eventually converges to a sparse set that maximizes the performance.
-Upon the convergence, the information can be preserved and transferred to the “non-zero" Gaussians.
-Another benefit over existing heuristic one-shot pruning approaches is that our removal of Gaussians does not rely on a one-time importance measurement but a long-term training process.
-As shown in Figure 2, existing one-shot, criterion-based methods lead to a sudden performance drop, followed by irreversible losses of 0.2 dB and 0.5 dB for importance-based and opacity-based methods, respectively, even with additional training iterations.
-In contrast, our GaussianSpa approach smoothly and optimally preserves knowledge, enabling more accurate view rendering.
-Moreover, the introduced variables, $\bm{z}$ and $\bm{\lambda}$, are considered constants, as we do not calculate their gradients.
-The “sparsifying" step is performed every tens iterations instead of every single iteration.
-Thus, the overall training speed is very similar to the original 3DGS.
-Algorithm 1 Procedure of “Optimizing-Sparsifying"
-Input: Gaussian opacity $\bm{a}$, 3DGS variables $\bm{\Theta}$, target number of Gaussians $\kappa$, penalty parameter $\delta$, feasibility tolerance $\epsilon$, maximum iterations $T$.
-Output: Optimized $\bm{a}$ and $\bm{\Theta}$.
-[…]
-2:$t$leftarrow00;
-3:while $\|\bm{a}-\bm{z}\|^{2}>\epsilon$ and $t\leq T$ do
-4:     Update $\bm{a}$ and $\bm{\Theta}$ via Eq. 14; $\triangleright$ “Optimizing" Step
-5:     Update $\bm{z}$ via Eq. 16; $\triangleright$ “Sparsifying" Step
-6:     Update $\bm{\lambda}$ via Eq. 17; $\triangleright$ Multiplier Update
-[…]
-8:end while
+respectively. Here, $\eta$ is the learning rate during training.
 Method Mip-NeRF 360 Tanks&Temples Deep Blending
@@
 “#G/M” represents number of million Gaussians.
+[…]
+In the “sparsifying" step, the minimization objective, Eq.
+15, targets only the auxiliary variable, $\bm{z}$, while the term of indicator function dominates Eq.
+15.
+This sub-problem essentially has a form of the proximal operator [40], $\mathrm{\bf{prox}}_{h}$, associated with the indicator function $h(\cdot)$.
+Thus, the solution of this sub-problem can be directly given by
+$\bm{z}\leftarrow\mathrm{\bf{prox}}_{h}(\bm{a}+\bm{\lambda}).$ (16)
+The proximal operator maps the opacity $\bm{a}$ to the auxiliary variable $\bm{z}$, which contains at most $\kappa$ non-zeros.
+This operator results in a projection solution [5], sparsifying $(\bm{a}+\bm{\lambda})$, by projecting it onto a set that satisfies $\|\bm{z}\|_{0}\leq\kappa$.
+More specifically, the projection keeps top-$\kappa$ elements and sets the rest to zeros.
+[…]
+$\bm{\lambda}\leftarrow\bm{\lambda}+\bm{a}-\bm{z}.$ (17)
+Overall, the “optimizing," “sparsifying," and multiplier update steps are performed alternately in the 3DGS training process until convergence, as summarized in Algorithm 1.
+Generally, we consider our solution convergences once the condition, $\|\bm{a}-\bm{z}\|^{2}\leq\epsilon$, is satisfied, or the iterations reach the predefined maximum number.
+Remark.
+The “optimizing" step optimizes over the original Gaussian variables and pushes the Gaussian opacity $\bm{a}$ close to the exactly sparse auxiliary variable $\bm{z}$ using a gradient-based approach, simultaneously satisfying the 3DGS performance and the sparsity requirements.
+The “sparsifying" step essentially prunes the auxiliary variable $\bm{z}$, which can be considered the exactly sparse version of $\bm{a}$.
+As those steps are performed alternately, summarized in Algorithm 1, the process eventually converges to a sparse set that maximizes the performance.
+Upon the convergence, most Gaussians become “zeros," and the information can be preserved and transferred to the “non-zero" Gaussians.
 3.4 Overall Workflow of GaussianSpa
@@
 The overall workflow of our GaussianSpa is illustrated in Figure 3.
-We have visualized the evolution of Gaussian opacity distribution along with the PSNR, trained on the Room scene, as shown in Figure 4.
-It is observed that the opacity distribution of GaussianSpa is significantly distinct from the original 3DGS – there is a clear gap between the “zero" Gaussians and the remaining Gaussians in GaussianSpa, while the PSNR keeps increasing.
+We visualize the evolution of Gaussian opacity distribution along with the PSNR, trained on the Room scene, as shown in Figure 5.
+The opacity distribution of GaussianSpa is significantly distinct from the original 3DGS – there is a clear gap between the “zero" Gaussians and the remaining Gaussians in GaussianSpa, while the PSNR keeps increasing.
 This means GaussianSpa has successfully imposed a substantial sparsity property and preserved information to the “non-zero" Gaussians.
@@
 4 Experiments
+Figure 6: Visualization for rendered Gaussian ellipsoids.
+(Add’l Figure 12 in the Supplementary Material).
+GaussianSpa renders high-frequency (e.g., air outlet of the train) and low-frequency (e.g., large-area sky) areas adaptively using appropriate numbers of Gaussians, respectively, more precisely matching the original textures, compared to the state-of-the-art Mini-Splatting [17] and the vanilla 3DGS.
 4.1 Experimental Settings
@@
 Our experimental server has two AMD EPYC 9254 CPUs and eight NVIDIA GTX 6000 Ada GPUs.
-We use the same checkpoint files as Mini-Splatting [20] for training, which are stored before our “optimizing-sparsifying” process to ensure fair comparisons.
+We use the same starting checkpoint files as Mini-Splatting [17] before our “optimizing-sparsifying” process to ensure fair comparisons.
 GaussianSpa starts “optimizing-sparsifying” at iteration 15K and removes “zero" Gaussians at iteration 25K.
+Figure 7: Quality-Reduction Rate curves on the (left) Kitchen and (right) Room scenes. GaussianSpa outperforms Mini-Splatting [17] at multiple reduction rates.
 4.2 Quantitative Results
@@
 These results quantitatively demonstrate the superiority of our GaussianSpa.
-We also plot quality curves with respect to the Gaussian reduction rate on the Room and Kitchen scenes in Figure 7 (more results are in Figure 13 in the Appendix), showing the robustness of GaussianSpa for different numbers of remaining Gaussians.
-Figure 7 shows that GaussianSpa outperforms the state-of-the-art method, Mini-Splatting [20], by an average of 0.5 dB across multiple Gaussian reduction rates.
-Figure 6: Visualization for rendered Gaussian ellipsoids.
-GaussianSpa adaptively reconstructs high-frequency (e.g., air outlet of the train, complex textures of the carpet) and low-frequency (e.g., large-area sky, simple textures of the carpet) regions using appropriate numbers of Gaussians, respectively, more precisely matching the original textures and shapes, compared to the state-of-the-art Mini-Splatting [20] and the vanilla 3DGS.
+We plot quality curves for the Gaussian reduction rate on the Room and Kitchen in Figure 7 (more results are in Figure 14 in the Supplementary Material), showing the robustness of GaussianSpa for different numbers of remaining Gaussians.
+GaussianSpa outperforms the state-of-the-art method, Mini-Splatting [17], by an average of 0.5 dB PSNR improvement across multiple Gaussian reduction rates.
+Figure 8: Visualized point clouds for the Room and Drjohnson scenes.
+With GaussianSpa, sparse Gaussians concentrate on high-frequency areas after our “optimizing-sparsifying"-integrated training process, ensuring the capability to capture detail-rich textures and shapes.
+On the contrary, the point clouds produced by other approaches cannot outline the contours.
 4.3 Visual Quality Results
-Figure 7: Quality-Reduction Rate curves on the (left) Kitchen and (right) Room scenes. GaussianSpa outperforms Mini-Splatting [20] across all reduction rates.
-As depicted in Figure 5, we present the rendered images on the Drjohnson and Counter scenes [5, 23] for visual quality evaluation, in comparison to LightGaussian [19], Mini-Splattin [20], and the vanilla 3DGS.
-Together with Figure 1, those outcomes demonstrate the superior rendering quality of our GaussianSpa on diverse scenes over existing approaches, consistent with the quantitative results in the previous subsection.
+Figure 4 shows the rendered images on the Drjohnson and Counter scenes [3, 20] for visual quality evaluation, in comparison to LightGaussian [16], Mini-Splatting [17], and the vanilla 3DGS.
+Together with Figure 1, these outcomes demonstrate GaussianSpa’s superior rendering quality on diverse scenes over existing approaches, consistent with the quantitative results in the previous subsection.
 For instance, as shown in Figure 4, GaussianSpa accurately renders the outlet behind the railings for the Playroom scene, while LightGaussian [16] and Mini-Splatting [17] lose this object.
 Furthermore, GaussianSpa fully recovers the original details of the zoom-in object in the Counter scene, while other methods outcome blurry surfaces.
-We comprehensively analyze the proposed GaussianSpa by visualizing the rendered Gaussian ellipsoids and final point clouds in Figure 6 and Figure 8, respectively.
+We analyze the proposed GaussianSpa by visualizing the rendered Gaussian ellipsoids and final point clouds in Figure 6 and Figure 8, respectively.
 Those visualized results demonstrate that GaussianSpa can adaptively control the density to match the details when sparsifying the Gaussians, ensuring high-quality sparse 3D representation.
-From Figure 6, we can observe that GaussianSpa uses fewer but larger Gaussians to represent the blue sky and uses dense Gaussians to render complex carpet textures, providing precise rendering of details in the zoom-in images.
-In contrast, Mini-Splatting [20] cannot fully reconstruct the details, leading to blurry rendered textures.
-Figure 8 further demonstrates our superiority in simplification among existing approaches.
-GaussianSpa adaptively reduce the Gaussians in the low-frequency areas and preserve more Gaussians in high-frequency areas that generally need more Gaussians to synthesize, resulting in sparse Gaussian point clouds that precisely outline the scene contours.
+From Figure 6 and Figure 12, we can observe that GaussianSpa uses fewer but larger Gaussians to represent the blue sky and uses dense Gaussians to render complex carpet textures, providing precise rendering of details in the zoom-in images.
+In contrast, Mini-Splatting [17] cannot fully present the details, leading to blurry rendered textures.
+Figure 8 further demonstrates the superiority of GaussianSpa in simplification over existing methods.
+GaussianSpa adaptively reduce the Gaussians in the low-frequency areas and preserve more Gaussians in high-frequency areas that generally need denser Gaussians to synthesize, resulting in sparse Gaussian point clouds that precisely outline the scene contours.
 5 Discussion
-5.1 Convergence Analysis
-Figure 8: Visualized point clouds for the Room and Drjohnson scenes.
-The numbers of remaining Gaussians in millions are displayed.
-With GaussianSpa, sparse Gaussians concentrate on high-frequency areas after our “optimizing-sparsifying"-integrated training process, ensuring the capability to capture detail-rich textures and shapes.
-On the contrary, the point clouds produced by other approaches cannot outline the contours.
-We analyze convergence behavior of the proposed GaussianSpa and appropriately set hyper-parameters, such as $\delta$, which controls the sparsity strength in Eq.
-10.
-As introduced in Section 3, GaussianSpa performs the “sparsifying" step every fixed number of iterations, referred to here as “interval" here.
-We plot loss curves for multiple $\delta$’s and intervals in Figure 9.
-It is seen that with different $\delta$ and interval values, the curves converge well at a similar rate during our “optimizing-sparsifying"-integrated training process.
-After removing “zero" Gaussians at iteration 25K, the loss curve still exhibits a consistent convergence, confirming the feasibility of GaussianSpa.
-Figure 9: Loss curves with multiple (left) $\delta$ and (right) interval settings.
-The curves show that GaussianSpa exhibits a good convergence behaviour in the “optimizing-sparsifying"-integrated training process.
 5.1 Generality of GaussianSpa
-GaussianSpa is a general Gaussian simplification framework.
-Existing importance measurement criteria can be applied to project the auxiliary variable $\bm{z}$ onto the sparse space in the “sparsifying" step, Eq.
-15, obtaining more compact 3DGS models with improved rendering quality.
-For example, we employ the criterion from LightGaussian [19] and plot the PSNR curves at multiple Gaussian reduction rates in Figure 13 in the Appendix.
+The proposed GaussianSpa is a general Gaussian simplification framework.
+Existing importance criteria can be applied to project the auxiliary variable $\bm{z}$ onto the sparse space in the “sparsifying" step, Eq.
+15, obtaining more compact 3DGS models with improved rendering performance.
+For example, we employ the criterion from LightGaussian [16] and plot the PSNR curves at multiple Gaussian reduction rates in Figure 14 in the Supplementary Material.
 It is observed that GaussianSpa consistently outperforms LightGaussian [16] with an average of 0.4 dB improvement for the Playroom and Kitchen scenes.
 5.2 Storage Compression
-While GaussianSpa focuses on simplifying and optimizing the number of Gaussians to obtain a sparse 3DGS representation, existing compression techniques such as SH optimization and vector quantization can be applied as add-ons to the simplified 3DGS model.
-As shown in Table 2, we present the final storage cost for the Mip-NeRF 360 dataset by incorporating the SH distillation and vector quantization methods proposed by LightGaussian [19].
-This comparison demonstrates that GaussianSpa can achieve the lowest storage consumption with additional compression techniques while maintaining superior rendering quality over existing hybrid compression methods, such as EfficientGS [31] and LightGaussian [19].
+While GaussianSpa focuses on simplifying and optimizing the number of Gaussians for a sparse 3DGS representation, existing compression techniques such as SH optimization and vector quantization can be applied as add-ons to the simplified 3DGS model.
+Table 2 (in Supplementary Material) shows the final storage cost for the Mip-NeRF 360 dataset by incorporating the SH distillation and vector quantization methods proposed by LightGaussian [16].
+This comparison demonstrates that GaussianSpa can achieve the lowest storage consumption while maintaining superior rendering quality over existing hybrid compression methods, such as EfficientGS [28] and LightGaussian [16].
 5.3 Relation to Other Optimization Algorithms
-The “optimizing-sparsifying" algorithm in our proposed GaussianSpa can be categorized into the class of optimization algorithms that project the target variable onto a constrained space, specifically the $\ell_{p}$ ball (where $p=0$ in our case).
-This approach is similar to foundational optimization algorithms in compressed sensing [16] and sparse learning [9], such as Ridge regression [24] and Lasso regression [48].
-These algorithms extend the least squares regression by incorporating $\ell_{p}$ regularization – Ridge regression adds the $\ell_{2}$ regularization, while the Lasso regression adds the $\ell_{1}$ regularization.
+The “optimizing-sparsifying" solution in our proposed GaussianSpa falls under optimization methods that project a target variable onto a constrained space, specifically the $\ell_{p}$ ball (where $p=0$ in our case).
+This approach is similar to foundational optimization algorithms in compressed sensing [13] and sparse learning [6], such as Ridge regression [21] with the $\ell_{2}$ regularization and Lasso regression [45] with the $\ell_{1}$ regularization.
 The regularization acts as a penalty, combined with quadratic loss, reformulating the problem as a convex optimization task.
-Method PSNR↑ SSIM↑ LPIPS↓ Storage↓
-EfficientGS [31] 27.38 0.817 0.216 98 MB
-LightGaussian [19] 27.28 0.805 0.243 42 MB
-GaussianSpa 27.85 0.825 0.214 25 MB
-Table 2: Storage comparison evaluated on the Mip-NeRF 360 dataset.
-GaussianSpa’s storage cost is reported based on the add-on compression methods (i.e., SH distillation and vector quantization) from LightGaussian [19].
-However, neither $\ell_{1}$ nor $\ell_{2}$ regularization yields truly sparse solutions; achieving genuine sparsity requires the $\ell_{0}$ regularization.
-Yet, when $p=0$, the problem, also known as the sparse coding problem [41], becomes nonconvex and is generally recognized as NP-hard [15, 37, 49].
-This complexity makes it computationally impractical when the dimension of the target variable is high.
-In such cases, Lasso (where $p=1$) offers a convex constraint region, serving as the tightest convex surrogate for the sparse coding problem, albeit with some sacrifice in exact sparsity [11].
-Another alternative solver suitable for this case is a greedy algorithm, such as the Matching Pursuit algorithm [33, 45].
-However, this algorithm cannot guarantee the global optima.
-Iterative thresholding (IT) is another direction to address this problem effectively [17, 14, 18].
-These algorithms perform a matrix multiplication and its adjoint at each iteration, followed by a scalar shrinkage step, which yields a simple yet highly effective structure.
-Furthermore, a closed-form solution for the global minimizer of the sparse coding problem can be obtained – even for non-convex cases – under the condition that a unitary matrix is involved [17].
+However, genuine sparsity requires $\ell_{0}$ regularization, which leads to a nonconvex, NP-hard problem [38, 12, 34, 46].
+Lasso (where $p=1$) provides the tightest convex approximation but sacrifices exact sparsity [8].
+Greedy algorithms like Matching Pursuit [30, 42] offer another approach, though they lack guaranteed global optima.
+Iterative thresholding (IT) [14, 11, 15] offers a different solution, using matrix multiplications and scalar shrinkage steps to yield a simple yet effective structure.
+Furthermore, a closed-form solution for the global minimizer can be obtained in cases with a unitary matrix [14], even for non-convex scenarios.
 Inspired by this, we introduce a duplicated variable (i.e., $\bm{z}$) and project it onto the $\ell_{0}$ ball, optimizing over two variables alternately, similar to the alternating direction method of multipliers [5].
-By leveraging the simplicity of the unitary case, the “sparsifying" step aligns with this scenario and thus admits an explicit solution.
-Hence, our algorithm simultaneously benefits from the sparsity imposed by the $\ell_{0}$ constraint and achieves fast convergence, guided by the explicit solution.
+Leveraging the unitary case’s simplicity, the “sparsifying" step aligns with this scenario and thus admits an explicit solution.
+Thus, our algorithm benefits from $\ell_{0}$-induced sparsity while achieving fast convergence guided by the explicit solution.
 6 Conclusion
-In this paper, we present an optimization-based 3DGS simplification framework, GaussianSpa, for compact and high-quality sparse view synthesis.
-GaussianSpa formulates 3DGS simplification as a constrained optimization problem with a sparsity constraint on the Gaussian opacities.
+In this paper, we present an optimization-based simplification framework, GaussianSpa, for compact and high-quality 3DGS.
+GaussianSpa formulates the simplification objective as a constrained optimization problem with a sparsity constraint on the Gaussian opacities.
 Then, we propose an “optimizing-sparsifying" solution to efficiently solve the formulated problem during the 3DGS training process.
-We comprehensively evaluate the proposed GaussianSpa on multiple datasets with quantitative results and qualitative analyses, demonstrating the superiority in rendering quality with fewer Gaussians compared to existing state-of-the-art approaches.
-Appendix
+Our comprehensive evaluations on multiple datasets with quantitative results and qualitative analyses demonstrate the superiority of GaussianSpa in rendering quality with fewer Gaussians compared to the state of the arts.
+Supplementary Material
 7 Additional Results
@@
 Those additional results demonstrate our GaussianSpa achieves stronger representational power for background and detail-rich areas such as walls, carpets, and ladders, showcasing superior rendering qualities.
-We also present additional point cloud views in Figure 12.
-This further illustrates that GaussianSpa creates a high-quality sparse 3D representation that adaptively uses more Gaussians to represent high-frequency areas.
-Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
-Drjohnson 3DGS 28.77 0.900 0.250 3.260
-Mini-Splatting 29.37 0.904 0.261 0.377
-GaussianSpa 29.89 0.913 0.243 0.450
-GaussianSpa 29.82 0.909 0.254 0.293
-Playroom 3DGS 30.07 0.900 0.250 2.290
-Mini-Splatting 30.72 0.914 0.248 0.417
-GaussianSpa 30.84 0.916 0.254 0.219
-GaussianSpa 30.84 0.916 0.254 0.219
-Average 3DGS 29.42 0.900 0.250 2.780
-Mini-Splatting 30.05 0.909 0.254 0.397
-GaussianSpa 30.37 0.914 0.249 0.335
-GaussianSpa 30.33 0.912 0.254 0.256
-Table 3: Deep Blending per scene results. 3DGS results are reported from [22]. Mini-Splatting [20] results are replicated using official code.
-Figure 10: Additional visual comparison on more scenes. The numbers of remaining Gaussians in millions are displayed.
-Figure 11: (Continue) Additional visual comparison on more scenes. The numbers of remaining Gaussians in millions are displayed.
-Figure 12: Additional visualized point clouds. The numbers of remaining Gaussians in millions are displayed.
-Figure 13: The first 13 sub-figures: Quality-#G (the number of Gaussians in millions) curves comparing GaussianSpa with Mini-Splatting [20] on multiple scenes.
-Our GaussianSpa consistently outperforms Mini-Splatting [20] with the same #G.
-The last two sub-figures: Quality-Reduction Rate curves comparing GaussianSpa with LightGaussian [19] on the Kitchen and Playroom scenes.
+Furthermore, Figure 12 and Figure 13 offer additional rendered Gaussian ellipsoids and point cloud views, respectively.
+These results further illustrate that GaussianSpa creates a high-quality sparse 3D representation that adaptively uses more Gaussians to represent high-frequency areas.
+8 Additional Discussion
+Method PSNR↑ SSIM↑ LPIPS↓ Storage↓
+EfficientGS [28] 27.38 0.817 0.216 98 MB
+LightGaussian [16] 27.28 0.805 0.243 42 MB
+GaussianSpa 27.85 0.825 0.214 25 MB
+Table 2: Storage comparison evaluated on the Mip-NeRF 360 dataset.
+GaussianSpa’s storage cost is reported based on the add-on compression methods (i.e., SH distillation and vector quantization) from LightGaussian [16].
+Convergence Analysis.
+We analyze the convergence behavior of GaussianSpa by examining the effects of hyper-parameters such as $\delta$, which controls the sparsity strength in Eq.
+10, and the interval at which the “sparsifying" step is performed, as described in Section 3.
+Figure 9 shows loss curves for various $\delta$ and interval values, indicating similar convergence rates during our “optimizing-sparsifying"-integrated training process.
+After removing “zero" Gaussians at iteration 25K, the loss curve continues to converge consistently, confirming GaussianSpa’s feasibility.
+Figure 9: Loss curves with multiple (left) $\delta$ and (right) interval settings.
+The curves show that GaussianSpa exhibits a good convergence behavior in the “optimizing-sparsifying"-integrated training process.
 Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
@@
 Table 3: MiP-NeRF360 per scene results. 3DGS results are reported from [19]. Mini-Splatting [17] results are replicated using official code.
+Figure 10: Additional visual comparison on more scenes. The numbers of remaining Gaussians in millions are displayed.
+Figure 11: (Continue) Additional visual comparison on more scenes. The numbers of remaining Gaussians in millions are displayed.
+Figure 12: Additional visualized Gaussian ellipsoids. The numbers of remaining Gaussians in millions are displayed.
+Figure 13: Additional visualized point clouds. The numbers of remaining Gaussians in millions are displayed.
+Figure 14: The first 13 sub-figures: Quality-#G (the number of Gaussians in millions) curves comparing GaussianSpa with Mini-Splatting [17] on multiple scenes.
+Our GaussianSpa consistently outperforms Mini-Splatting [17] with the same #G.
+The last two sub-figures: Quality-Reduction Rate curves comparing GaussianSpa with LightGaussian [16] on the Kitchen and Playroom scenes.
 Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
@@
 GaussianSpa 23.98 0.852 0.180 0.269
-Table 5: Tanks&Temples per scene results. 3DGS results are reported from [22]. Mini-Splatting [20] results are replicated using official code.
+Table 4: Tanks&Temples per scene results.
+Scene Method PSNR↑ SSIM↑ LPIPS↓ # G (M)↓
+Drjohnson 3DGS 28.77 0.900 0.250 3.260
+Mini-Splatting 29.37 0.904 0.261 0.377
+GaussianSpa 29.89 0.913 0.243 0.450
+GaussianSpa 29.82 0.909 0.254 0.293
+Playroom 3DGS 30.07 0.900 0.250 2.290
+Mini-Splatting 30.72 0.914 0.248 0.417
+GaussianSpa 30.84 0.916 0.254 0.219
+GaussianSpa 30.84 0.916 0.254 0.219
+Average 3DGS 29.42 0.900 0.250 2.780
+Mini-Splatting 30.05 0.909 0.254 0.397
+GaussianSpa 30.37 0.914 0.249 0.335
+GaussianSpa 30.33 0.912 0.254 0.256
+Table 5: Deep Blending per scene results.